Lou wants to make a coffee mixture to sell. He is going to mix Sumatra coffee which cost $2.50 per pound with Colombian coffee which cost $3.75 per pound. He wants to make 50 pounds of mix and he wants the cost of the mix to be $3.35 per pound. How many pounds of each will he need?
Answer by ankor@dixie-net.com(22740)

Lou wants to make a coffee mixture to sell.
He is going to mix Sumatra coffee which cost $2.50 per pound with Colombian coffee which cost $3.75 per pound.
He wants to make 50 pounds of mix and he wants the cost of the mix to be $3.35 per pound.
How many pounds of each will he need?
:
let s = amt of Sumatra coffee required
let c = amt of Columbian coffee
:
Two equations
The resulting amt is to be 50 lb, therefore
s + c = 50
s = (50-s), use this form for substitution
:
and the mixture equation
2.50s + 3.75c = 3.35(50)
replace s with (50-c)
2.50(50-c) + 3.75c = 167.5
125 - 2.50c + 3.75c = 167.5
-2.50c + 3.75c = 167.5 - 125
1.25c = 42.5
c = 42.5/1.25
c = 34 lb of Columbian coffee
and
50 - 34 = 16 lb of Sumatra coffee
:
It's up to you to check the math here